Pleanse help me, I am mom trying to help my daughter with this problem but I can't get the right answer.
A website allows users to download individual songs or an entire album. All individual songs cost the same to download. Ryan pays $14.94 to download 5 individual songs and 1 album. Seth pays $22.95 to download 3 individual songs and 2 albums. How much does the website charge to download a song? an entire album?

Answer by checkley79(3341)

5S+1A=14.45 MULTIPLY THIS EQUATION BY -2 & ADD.
3S+2A=22.95
-10S-2A=-28.90
-----------------
-7S=-5.95
S=-5.95/-7
5S=.85 ANS. FOR THE PRICE OF A SINGLE.
5*.85+A=14.45
4.25+A=14.45
A=14.45-4.25
A=10.20 ANS. FOR THE COST OF A SINGLE.
PROOFl:
3*.85+2*10.20=22.95
2.55+20.40=22.95
22.95=22.95